# Data-Augmented Deep Learning for Downhole Depth Sensing and Field Validation

### v2

@@ III-C Results and Analysis
To explore the validity and importance of augmentation methods, several real and complete waveforms are experimented to verify the models’ ability of predicting collar positions. The experimental results are indexed and divided into 3 groups, as illustrated in Fig. 4, with curve indexes corresponding to those in Tables I and II. The first group (Cfgs. 1-6) examines normalization methods, label distribution methods, and cropping methods. The second group (Cfgs. 7-12) investigates label smoothing, noise injection, amplitude jittering, and time scaling. The final group (Cfgs. 13-18) explores optimal combination configurations.
@@ III-C1 Fundamental Preprocessing Requirements
Similar phenomena occur with min-max normalization versus standardization (corresponding to Cfgs. 3-5), and fixed cropping versus random cropping (corresponding to Cfgs. 2 and 3), as illustrated in Figs. 5(b,c). Waveforms processed with min-max normalization prove challenging for model training. Furthermore, models using fixed cropping learn only sliding window positions rather than waveform characteristics.
@@ III-C2 Generalization Enhancement Methods
To investigate the effects of LSR, noise injection, amplitude jittering, and time scaling, Cfgs. 7-10 based on Cfg. 3 are experimented. Comparison of curves in Figs. 4(d-i) reveals that: (a) the final CE loss for Cfgs. 8-10 is lower than that for control Cfg. 3, while Cfg. 7 shows higher loss; (b) the classifier F1 scores of Cfgs. 7-9 are similar, while Cfg. 10 shows slightly lower performance; (c) LSR convergence is slower than others. Performance comparison between Cfgs. 7-10 and control Cfg. 3, as tableted in Table II, demonstrates that: (a) LSR exhibits higher CE loss but superior F1 score in waveform evaluation, with 0.024 improvement on the moderate inference waveform; (b) although amplitude jittering and time scaling show lower F1 scores on validation set evaluation is compared to the control configuration, their waveform evaluation F1 scores are higher, with improvements of 0.012 and 0.037 respectively on the moderate inference waveform. These results suggest that LSR enhances generalization capability at the cost of convergence speed, while amplitude jittering and time scaling also improve generalization. However, noise injection performs worse than all other configurations. Additional experiments reveal that small noise provides limited benefits to performance, while large noise impairs performance. This likely occurs because real-world waveforms inherently contain small noise, and additional large noise hinders model learning waveform characteristics.
To investigate the effects of multiple sampling, Cfgs. 11-12 based on Cfg. 3 are experimented. Configurations with multiple sampling converge substantially faster than the control configuration, as shown in Figs. 4(d-i), because multiple sampling increases iterations per epoch. The F1 score improves up to 0.045 in evaluation. In summary, LSR, amplitude jittering, time scaling, and multiple sampling significantly enhance generalization capability, while noise injection provides limited benefits.
@@ III-C3 Optimal Configuration Identification
Based on these conclusions, Cfgs. 13-18 are experimented to identify optimal combinations while comparing with Cfgs. 3, 6 and 7, as illustrated in Figs. 4(j-o) and  5(c-e) and tabulated in Table II. Several noteworthy phenomena emerge from these experiments.
First, metrics of configurations with 100x multiple sampling fluctuate more dramatically than those with 20x, suggesting the need for smaller initial learning rates when applying extensive multiple sampling. Second, simultaneous use of amplitude jittering and time scaling unexpectedly reduces the performance. We hypothesize that amplitude scaling compromises standardization benefits while improving generalization, as the relative amplitude of waveforms contains critical information. Third, Cfg. 13 achieves the highest CE loss and validation set F1 score among all TAN model configurations, yet its F1 score on the moderate interference waveform does not show corresponding improvement, as illustrated in Figs. 4(m-o). This occurs because multiple sampling introduces more randomly preprocessed samples, enhancing generalization capability.
@@ III-C4 Performance Validation and Key Findings
These findings establish that the standardization, LDS, and random cropping are fundamental preprocessing requirements for collar recognition model training, while LSR, time scaling, and multiple sampling significantly enhance model generalization capability. The data augmentation methods achieve F1 score improvements of up to 0.063 for TAN model waveform evaluation and up to 0.048 for MAN model evaluation, compared to configurations using only fundamental methods, representing notable performance enhancements.
IV CONCLUSIONS
This work developed the SCV to sample CCL signals downhole in the wellborn for datasets construction. We systematically analyzed the contributions of various data augmentation methods using benchmark evaluations on our proposed neural network recognition methodology. Experimental validation with authentic CCL waveforms confirm the effectiveness and validity of proposed approaches. This research addresses the existing gaps in data augmentation methodologies for training casing collar recognition models in CCL data-limited environments.
ACKNOWLEDGMENT
This work is supported by NSFC under project No. 62404033 and 62404034. This work is also supported by China National Petroleum Corporation Logging Co., Ltd. (CNLC) under project No. CNLC2023-7A01.

### v3

@@ III-B Training and Validation
The models are trained with a batch size of 16 for 100 epochs utilizing cross-entropy loss (CE loss) with an adaptive moment estimation (Adam) optimizer. Key configurations (Cfgs.) are presented in Tables I and II. The training progress and the optimal model performance for each configuration are illustrated in Fig. 4. Additionally, model inference and post-processing were evaluated on two full-length waveforms characterized by mild and moderate interference, respectively. The mild interference waveform contains 52 collars with maximum depth of $512.2\text{\,}\mathrm{m}$, while the moderate interference waveform contains 77 collars with maximum depth of $771.1\text{\,}\mathrm{m}$. Crucially, the waveform fragments from these evaluation waveforms were excluded from the training set. The precision, recall, and F1 score results are tabulated in Tables I and II, with the aggregated results for both waveforms are tabulated in Table III.
III-C Results and Analysis
TABLE III: Comparison of Performance with Other Works
Network Performance Method
Tests Acc P R F1
Cfg. 3 129 0.920 0.940 0.977 0.958 CCL + TAN + LDS
Cfg. 6 129 0.946 1 0.946 0.972 CCL + MAN + LDS
Cfg. 13 129 0.970 1 0.970 0.985 CCL + TAN + LDS + Data Augmentation
Cfg. 14 129 0.992 1 0.992 0.996 CCL + MAN + LDS + Data Augmentation
[11] 269 0.974 1 0.942 0.970 CCL + CNN
[11] 269 0.948 1 0.884 0.939 CCL + LSTM
[11] 269 0.978 0.959 0.991 0.975 CCL + CNN-LSTM
[32] 579 0.973 0.988 0.985 0.986 CCL + Dynamic amplitude threshold + Physical plausibility
[29] 8 1 CCL + Relative amplitude
[19] CCL + Cross correlation + Predifined threshold
[17] CCL + Wavelet transform
“–” indicates “Not mentioned”; “P”, “R” are precision and recall, respectively.
To explore the validity and importance of augmentation methods, several real and complete waveforms are experimented to verify the models’ ability of predicting collar positions. The experimental results are indexed and divided into 3 groups, as illustrated in Fig. 4, with curve indexes corresponding to those in Tables I and II. The first group (Cfgs. 1–6) examines normalization methods, label distribution methods, and cropping methods. The second group (Cfgs. 7–12) investigates label smoothing, noise injection, amplitude jittering, and time scaling. The final group (Cfgs. 13–18) explores optimal combination configurations.
@@ III-C1 Fundamental Preprocessing Requirements
Similar phenomena occur with min-max normalization versus standardization (corresponding to Cfgs. 3–5), and fixed cropping versus random cropping (corresponding to Cfgs. 2 and 3), as illustrated in Figs. 5(b,c). Waveforms processed with min-max normalization prove challenging for model training. Furthermore, models using fixed cropping learn only sliding window positions rather than waveform characteristics.
@@ III-C2 Generalization Enhancement Methods
To investigate the effects of LSR, noise injection, amplitude jittering, and time scaling, Cfgs. 7–10 based on Cfg. 3 are experimented. Comparison of curves in Figs. 4(d–i) reveals that: (a) the final CE loss for Cfgs. 8–10 is lower than that for control Cfg. 3, while Cfg. 7 shows higher loss; (b) the classifier F1 scores of Cfgs. 7–9 are similar, while Cfg. 10 shows slightly lower performance; (c) LSR convergence is slower than others. Performance comparison between Cfgs. 7–10 and control Cfg. 3, as tableted in Table II, demonstrates that: (a) LSR exhibits higher CE loss but superior F1 score in waveform evaluation, with 0.024 improvement on the moderate inference waveform; (b) although amplitude jittering and time scaling show lower F1 scores on validation set evaluation is compared to the control configuration, their waveform evaluation F1 scores are higher, with improvements of 0.012 and 0.037 respectively on the moderate inference waveform. These results suggest that LSR enhances generalization capability at the cost of convergence speed, while amplitude jittering and time scaling also improve generalization. However, noise injection performs worse than all other configurations. Additional experiments reveal that small noise provides limited benefits to performance, while large noise impairs performance. This likely occurs because real-world waveforms inherently contain small noise, and additional large noise hinders model learning waveform characteristics.
To investigate the effects of multiple sampling, Cfgs. 11–12 based on Cfg. 3 are experimented. Configurations with multiple sampling converge substantially faster than the control configuration, as shown in Figs. 4(d–i), because multiple sampling increases iterations per epoch. The F1 score improves up to 0.045 in evaluation.
In summary, LSR, amplitude jittering, time scaling, and multiple sampling significantly enhance generalization capability, while noise injection provides limited benefits. By comparing Cfg. 13 and Cfg. 14 with Cfg. 3 and Cfg. 6, respectively, these data augmentation methods achieve F1 score improvements of up to 0.027 for TAN model and 0.024 for MAN model.
@@ III-C3 Optimal Configuration Identification
Based on these conclusions, Cfgs. 13–18 are experimented to identify optimal combinations while comparing with Cfgs. 3, 6 and 7, as illustrated in Figs. 4(j–o) and  5(c–e) and tabulated in Table II. Several noteworthy phenomena emerge from these experiments.
First, metrics of configurations with $100\times$ multiple sampling fluctuate more dramatically than those with $20\times$, suggesting the need for smaller initial learning rates when applying extensive multiple sampling. Second, simultaneous use of amplitude jittering and time scaling unexpectedly reduces the performance. We hypothesize that amplitude scaling compromises standardization benefits while improving generalization, as the relative amplitude of waveforms contains critical information. Third, Cfg. 13 achieves the highest CE loss and validation set F1 score among all TAN model configurations, yet its F1 score on the moderate interference waveform does not show corresponding improvement, as illustrated in Figs. 4(m–o). This occurs because multiple sampling introduces more randomly preprocessed samples, enhancing generalization capability.
Notably, both full-length evaluation waveforms contain several dozens of collar signatures, effectively functioning as a small-scale evaluation set. Consequently, it is plausible to achieve near-perfect or perfect results (an F1 score of 1.0) provided the model possesses sufficient accuracy. Furthermore, in the case of waveforms with mild interference, the collar signatures exhibit greater separability, thereby facilitating the attainment of perfect identification performance.
@@ III-C4 Performance Validation and Key Findings
These findings establish that the standardization, LDS, and random cropping are fundamental preprocessing requirements for collar recognition model training, while LSR, time scaling, and multiple sampling significantly enhance model generalization capability. The data augmentation methods achieve F1 score improvements of up to 0.063 for TAN model waveform evaluation and up to 0.048 for MAN model evaluation, compared to configurations using only fundamental methods, representing notable performance enhancements. Furthermore, the baseline models TAN and MAN exhibit performance comparable to prior works, as tabulated in Table III. However, applying the proposed data augmentation methods yields F1 score improvements of up to 0.045 for TAN model and 0.057 for MAN model relative to other approaches. This observation demonstrates that the proposed data augmentation methods are effective even when the underlying model architecture is not optimal.
IV Conclusion
This work developed the SCV to acquire raw downhole CCL signals for dataset construction. The contributions of various data augmentation methods were systematically analyzed through baseline evaluations using the proposed neural network recognition methodology. Experimental validation with field CCL logs confirms the effectiveness and validity of proposed approaches. Results indicate that standardization, LDS, and random cropping are fundamental preprocessing prerequisites for training collar recognition models, while LSR, time scaling, and multiple sampling significantly enhance model generalization capabilities. Incorporating the proposed augmentation methods into the two baseline models results in maximum F1 score improvements of 0.027 and 0.024 for the TAN and MAN models, respectively. Furthermore, applying these techniques yields F1 score gains of up to 0.045 for the TAN model and 0.057 for the MAN model compared to prior studies. This research addresses the existing gaps in data augmentation methodologies for training casing collar recognition models under CCL data-limited conditions, and provides a technical foundation for the future automation of downhole operations.
Biographies
[Uncaptioned image]     Si-Yu Xiao received the B.S. degree in Microelectronic Science and Engineering from University of Electronic Science and Technology of China (UESTC) in 2018. He is currently pursuing the Ph.D. degree in Electronic Science and Technology. His research interests include digital circuit design, edge computing, artificial intelligence and their applications.
[Uncaptioned image]   Xin-Di Zhao received the Bachelor of Engineering degree in Software Engineering from Southwest Petroleum University. He is now a Level 2 Engineer in perforation technology research at the Southwest Branch of China National Petroleum Corporation Logging Co., Ltd. His current research includes new perforation tools, technologies and methods.
[Uncaptioned image]   Tian-Hao Mao is currently pursuing a Bachelor’s degree in Integrated Circuit Design and Integrated Systems at the University of Electronic Science and Technology of China. His academic interests focus on digital circuit design, system-on-chip (SoC) architecture, and algorithm optimization for hardware acceleration.
[Uncaptioned image]   Yi-Wei Wang received the B.S. degree in microelectronics from the University of Electronic Science and Technology of China, Chengdu, China. He is currently working toward the M.S. degree with the School of Integrated Circuits, University of Electronic Science and Technology of China. His research interests include artificial intelligence algorithms, neural network accelerator design, and their applications in artificial intelligence.
[Uncaptioned image]   Yu-Qiao Chen is currently pursuing the M.S. degree in Integrated Circuit Science and Engineering at the University of Electronic Science and Technology of China, where he also received the B.S. degree in Microelectronic Science and Engineering. His research interests cover digital circuit design, System-on-Chip (SoC), and neural network algorithms.
[Uncaptioned image]   Hong-Yun Zhang received the Master of Science degree in Communication Networks and Software from the University of Surrey, UK. He is now a Level 4 Engineer in perforation technology research at the Southwest Branch of China National Petroleum Corporation Logging Co., Ltd. His current research includes new perforation tools, technologies and methods.
[Uncaptioned image]   Jian Wang received the Bachelor of Engineering degree in Software Engineering from Chengdu University of Technology. He is now a Level 4 Engineer in perforation technology research at the Southwest Branch of China National Petroleum Corporation Logging Co., Ltd. His current research includes new perforation tools, technologies and methods.
[Uncaptioned image]   Jun-Jie Wang received the Ph.D. degree in microelectronics from the University of Electronic Science and Technology of China. He is now a researcher at the University of Electronic Science and Technology of China. His current research interests include digital circuit design, nonvolatile memory devices, and their applications in artificial intelligence.
[Uncaptioned image]   Shuang Liu received the Ph.D. degree in microelectronics from the University of Electronic Science and Technology of China. He is now a postdoctoral researcher at the University of Electronic Science and Technology of China. His current research interests include processing-in-memory circuits and neuromorphic systems.
[Uncaptioned image]   Tu-Pei Chen is a tenured faculty member in the School of Electrical and Electronic Engineering at Nanyang Technological University, Singapore, where he has been teaching and conducting research in the field of Microelectronics for over 25 years. His current research interests include on-chip ESD and latch-up protection, memory devices, memory-based computing (in-memory computing, neuromorphic computing), and thin-film transistors and applications.
[Uncaptioned image]   Yang Liu received the B.Sc. degree in microelectronics from Jilin University, China, in 1998 and the Ph.D. degree from Nanyang Technological University, Singapore, in 2005. From May 2005 to July 2006, he was a Research Fellow with Nanyang Technological University, Singapore. In 2006, he was awarded the prestigious Singapore Millennium Foundation Fellowship. In 2008, he joined the School of Microelectronics, University of Electronic Science and Technology, China, as a full professor. He is the author or coauthor of over 130 peer-reviewed journal papers and more than 100 conference papers. He has been awarded one US patent and more than 30 China patents also. His current research includes memristor neural network system, neuromorphic computing ICs, and AI-RFICs.